OpenStudy (volleyballlover55):

The Grand Canyon is approximately 2^9 kilometers long. Mariner Valley is a canyon on Mars that is approximately 2^12 kilometers long. About how many times longer is Mariner Valley than the Grand Canyon? 2 2^3 2^21 2^108

OpenStudy (volleyballlover55):

@Directrix

OpenStudy (volleyballlover55):

@satellite73

Directrix (directrix):

Grand Canyon 2^9 kilometers Mariner Valley 2^12 kilometers MV ---- = 2^12 ==== = 2 ^ (12 - 9) = 2 to the what power times as large GC 2^9 @volleyballlover55

OpenStudy (volleyballlover55):

is it 2^108?

Directrix (directrix):

2 ^ (12 - 9) What is 12 - 9 =

OpenStudy (volleyballlover55):

3

Directrix (directrix):

2^3

OpenStudy (volleyballlover55):

2^3 which is B?

Directrix (directrix):

Yes. MV is 2^3 or 8 times as long as the GC.

OpenStudy (volleyballlover55):

ok :)